OpenStudy (anonymous):

Given the function g(x) = -4x + 5, compare and contrast g(2) and g(-4). Choose the statement that is true concerning these two values. Answer The values of g(2) and g(-4) cannot be compared. The value of g(2) is the same as the value of g(-4). The value of g(2) is smaller than the value of g(-4). The value of g(2) is larger than the value of g(-4).

OpenStudy (anonymous):

did you plug in the x=2 and x=-4? where ever you see an x, replace that with 2 or -4. for example: f(2) = -4(2) + 5 = -8 + 5 = -3 now you do that for x=-4, f(-4) =